OpenStudy (anonymous):

what is the volume of a sphere with a diameter of 15 cm?

OpenStudy (bahrom7893):

d = 15 r = 7.5 V = (4/3)*pi*(7.5)^3 = 1767 cm

OpenStudy (bahrom7893):

check my arithmetic..

OpenStudy (anonymous):

he is right

OpenStudy (anonymous):

done the math is like you

OpenStudy (anonymous):

sorry to keep bothering yall with the same types of ?'s but i jus cant remember them an im only triina get thru this quiz

OpenStudy (anonymous):

just do like this mate,diameter is equal to 2r then as bahrom said we find r=15/2

OpenStudy (anonymous):

volume is like this V=\[4piR^{3}/3 \]

OpenStudy (sgadi):

Hi Bahrom, Your work is correct (Except the units). The answer is \[\approx 1767.09 cm^3\]